OpenStudy (anonymous):

Verify the equation: cos(4x) = 1 - 6sin^2(x) + 4sin^4(x) The answer I came up with is cos(4x) = 1 - 8sin^2(x) + 8sin^4(x) What am I doing wrong?

OpenStudy (anonymous):

expansion of terms

OpenStudy (anonymous):

Use half angle formula.

OpenStudy (anonymous):

\[ \sin^2(\theta) = \frac{1-\cos(2\theta)}{2} \]

OpenStudy (anonymous):

Oh wait, you're using double angle? \[ \cos(2\theta) = \cos^2(\theta)-\sin^2(\theta) \]

OpenStudy (anonymous):

yes ... double angles

OpenStudy (anonymous):

You're probably just messing up your algebra.

OpenStudy (anonymous):

You're probably right :) ... but I've reworked it several times

OpenStudy (anonymous):

If I type out the work, will you take a look at it and see where I messed up?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

cos(4x) = cos(2x + 2x) = cos(2x)cos(2x) - sin(2x)sin(2x) = [(cos^2(x) - sin^2(x)] [(cos^2(x) - sin^2(x)] - sin(2x)sin(2x) = [1 - 2sin^2(x)][1 - 2sin^2(x)] - sin(2x)sin(2x) = [1 - 4sin^2(x) + 4sin^4(x) - [2sin(x)cos(x) * 2sin(x)cos(x)] = [1 - 4sin^2(x) + 4sin^4(x) - 4sin^2(x)cos^2(x) = [1 - 4sin^2(x) + 4sin^4(x) - [4sin^2(x) * (1 - sin^2(x)] = [1 - 4sin^2(x) + 4sin^4(x) - [4sin^2(x) - 4sin^4(x)] = 1 - 8sin^2(x) + 8sin^4(x) Whew ... :)

OpenStudy (anonymous):

Looking over it now...

OpenStudy (anonymous):

= [(cos^2(x) - sin^2(x)] [(cos^2(x) - sin^2(x)] - sin(2x)sin(2x) = [1 - 2sin^2(x)][1 - 2sin^2(x)] - sin(2x)sin(2x) This step doesn't make sense to me.

OpenStudy (anonymous):

Wait nevermind, I get it now.

OpenStudy (anonymous):

Seems that your work is correct.

OpenStudy (anonymous):

:) Thank you so much for checking it for me! I really appreciate it!